OpenStudy (anonymous):

-92+(-7) DO YOU TAKE THE -92 AND SUBTRACT IT FROM -7

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

K

OpenStudy (alexandervonhumboldt2):

no you -92+(-7)=-92-7

OpenStudy (anonymous):

It's literally the same as -92-7.

OpenStudy (alexandervonhumboldt2):

-92-7=-99

OpenStudy (igreen):

Subtract -7 from -92 is the same as: -92 - (-7) Which is incorrect..

OpenStudy (igreen):

Subtracting*

OpenStudy (alexandervonhumboldt2):

yeah